OpenStudy (anonymous):

The pool at a park is circular. You want to find the equation of the circle that is the boundary of the pool. Find the equation if the area of the pool is 400 square feet and (0, 0) represents the center of the pool.

OpenStudy (anonymous):

flutter off

OpenStudy (anonymous):

Hi there! I would like to help you! Would you like to learn the material with me?

OpenStudy (anonymous):

area = pi * r^2 r = Sqrt[area/Pi] the equation of the circle is x^2+y^2=r^2 x^2+y^2=400/Pi